OpenStudy (logan13):

The measure of an angle is 51°. What is the measure of its supplement? A. 39° B. 49° C. 129° D. 139°

rishavraj (rishavraj):

supplementary angle adds upto 180 :)) so x + 51 = 180

OpenStudy (theprofessor35):

So, a supplement has a total of 180

OpenStudy (theprofessor35):

Solve for x 180 - 51 = x

OpenStudy (alex6799):

180-51=129

OpenStudy (theprofessor35):

x = 129 the value of the other angle is 129

OpenStudy (theprofessor35):

Done! Simple?